OpenStudy (anonymous):

Solve this system of equations using substitution. Put your answer in ordered pair form, (x, y). x - 2y = -4 2x - y = 1

OpenStudy (agentjamesbond007):

x=(2y-4) plug to other equation 2(2y-4)-y=1 4y-8-y=1 combine like terms 3y=9 solve for y y=3 that's your y-coordinate plug into any equation for x plugging into first x - 2(3) = -4 x -6 = -4 x = 2 so the solution is (2,3)

OpenStudy (anonymous):

thanks a million, you just saved my @$$

OpenStudy (agentjamesbond007):

You're Welcome! Glad to help.